OpenStudy (anonymous):

Solve for x. Round your answer to two decimal places. Show your work for full credit.

OpenStudy (anonymous):



OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

@majdishokri

OpenStudy (dumbcow):

SOH CAH TOA \[\sin 32 = \frac{20}{x}\]

OpenStudy (anonymous):

which is a,b, and c? @MeowMeow123

OpenStudy (anonymous):

you have two way each other true but the best like dumbcow way

OpenStudy (anonymous):

C is the one you're trying to find, and i don't think the a and b order matters

OpenStudy (dumbcow):

@MeowMeow123 , 32 is the angle not the 3rd side

OpenStudy (anonymous):

\[32^{2}+20^{2}=x ^{2} \to x=\sqrt{32^{2}+20^{2}}\]

OpenStudy (anonymous):

37.73

OpenStudy (anonymous):

@majdishokri

OpenStudy (dumbcow):

@ilyautumn , am i wrong this is a simple pythagorean thm problem ?

OpenStudy (anonymous):

you're right .

OpenStudy (anonymous):

i'm just confused.

OpenStudy (anonymous):

@dumbcow but you can solve like this i think your way ??

OpenStudy (dumbcow):

what a weird problem it actually turns out the 3rd side is basically 32 as well ....haha :)

OpenStudy (dumbcow):

\[\tan(32) = \frac{20}{y}\] \[y = \frac{20}{\tan(32)} = 32.007\]

OpenStudy (dumbcow):

@ilyautumn \[\sin(32) = \frac{20}{x}\] \[x = \frac{20}{\sin(32)} = 37.742\]

OpenStudy (anonymous):

thats pretty close to the answer that i got when i did it the other way